Calculate the fair present values of the following bonds, all of which pay interest semiannually, have a face value of $1,000, have 8 years remaining to maturity, and have a required rate of return of 13 percent. a. The bond has a 8.4 percent coupon rate. (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16)) b. The bond has a 10.4 percent coupon rate. (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16) c. The bond has a 13 percent coupon rate. (Do not round intermediate calculations.) a. Fair present value b. Fair present value c. Fair present value
Answer #1

please find below the solution..please provide rating...

a) we have to use financial calculator to sovle this problem
Put in calculator
FV 1,000
PMT =1000*8.4%/2 42
I =13%/2 6.50%
N =8*2 16
compute PV ($775.34)
Therefore fair present value = $775.34
b) we have to use financial calculator to sovle this problem
Put in calculator
FV 1,000
PMT =1000*10.4%/2 52
I =13%/2 6.50%
N =8*2 16
compute PV ($873.02)
Therefore fair present value = $873.02
c) we have to use financial calculator to sovle this problem
Put in calculator
FV 1,000
PMT =1000*13%/2 65
I =13%/2 6.50%
N =8*2 16
compute PV ($1,000.00)
Therefore fair present value = $1,000.00